A city plans to plant grass seed in a public playground that has the shape of a triangle with a height of H = 60 m and a base of B = 20 m. Each bag of grass seed will seed
120 m2.
How many bags of seed should be purchased?

Answer by Boreal(15235)

A=(1/2)bh=(1/2)*20*60=600 m^2
600m^2/120 m^2/bag= 5 bags ANSWER